Find the inverse of the given one to one function f(x) = (x-1)/(x+2) Give the domain and the range of f and of inverse
Answer by stanbon(75887)

Find the inverse of the given one to one function f(x) = (x-1)/(x+2) Give the domain and the range of f and of inverse
---------
Domain of f(x): All Real Numbers except x = -2
Range of f(x): All Real Numbers except y = x/x = 1
-----------------
Inverse ??
1st: Interchange x and y:
x = (y-1)/(y+2)
-----
xy+x = y-1
-----
Solve for "y":
y-xy = x+1
----
y(1-x) = x+1
---
y = (x+1)/(1-x)
---------------------------
Domain of f^-1(x)::: All Real Numbers except x = 1
Range of f^-1(x):::: All Real Numbers except y = 1
====================================
Cheers,
Stan H.